OpenStudy (anonymous):

if 3sinA=1,find sin3A

OpenStudy (lucaz):

I don't know how to solve only with the variable A, but if you're allowed to use calculator you see that \[\Large{3\times sin(A)=1~then:~sin(A)=\frac{1}{3}}\] you can find the arcsine of 1/3 and subtitute it by A and find sin(3A) \[\large{arcsine(\frac{1}{3})=19.47~approx.~then:~sin(3\times19.47)=sin(58.41)}\]

OpenStudy (lucaz):

almost sin(60) wich would be \[\large{\frac{\sqrt{3}}{2}}\]